OpenStudy (anonymous):

A child is creating a pyramid with building blocks. The top three levels include 4 blocks, 10 blocks, and 16 blocks. How many blocks would be needed for a pyramid 25 levels tall? Use complete sentences to explain how a sum of an arithmetic series was applied?

OpenStudy (ash2326):

a1=4 a2=10 a3=16 an=a1+(n-1)d here the common difference d=a2-a1=6 for 25 level, let's find the number of blocks n=25 a25=a1+(25-1)*d a25=4+24*6 a25=148 to find the total no. of blocks we have to sum the number of blocks used in each level. Sum of a arithmetic progression is given by=(n/2)*(a1+an) (25/2)(a1+a25) n=25 a1=4 a25=148 =25*(4+148)/2 25*152/2 Sum=1900 blocks for the pyramid

OpenStudy (anonymous):

Thanks So Much :)